a chemist has 2 solutions, one has 40% concentration the other has 25%. how many liters will be mixed of each to get 113 liters of a 26% concentration
Answer by rfer(16322)

.4x+.25(113-x)=.26*113
.4x+28.25-.25x=29.38
.15x=1.13
x=7.53 lt of 40%
113-x=105.47 lt of 25%